y C G I C+G+I C+G+I(2)
0 250 100 150 500 550
500 650 100 150 900 950
1000 1050 100 150 1300 1350
1500 1450 100 150 1700 1750
2000 1850 100 150 2100 2150
2500 2250 100 150 2500 2550
3000 2650 100 150 2900 2950
3500 3050 100 150 3300 3350
4000 3450 100 150 3700 3750
4500 3850 100 150 4100 4150

1. The consumption function is C=250+0.8y. How much are a (the intercept) and MPC (Marginal Propensity to Consume)?

2. What is the value of the spending multiplier?

3. How much is equilibrium y?

4. Reduce G to zero. Fill in the new C+G+I(2) (was already done in chart). By how much has equilibrium y changed after G decreased?
Answer #1
  1. The consumption function is C=250+0.8y. How much are a (the intercept) and MPC (Marginal Propensity to Consume)= 0.8

Intercept is =$250

2.

MPC=0.8

MPS=1-MPC

=1-0.8

=0.2

Spending multiplier=1/MPS

=1/0.2

=5

3.

Since equilibrium determined by the following condition

Y=C+I+G

Hence equilibrium output (Y) will be $2,500

4.

If G reduces to zero, then New equilibrium will be determined by the following condition

Y=(C+I+G)2

Hence equilibrium (Y) will be 2,000.

New Table.

C 500 100 100 100 G 250 650 6501 1050 1450 1850 22501 2650 30501 3450 3850 100 1000 1500 2000 2500 30001 3500 4000 4500 C+G+1